Answer all after part A

An RLC series circuit has a 2.25 Ohm resistor, a 1
Answer #1

given

R = 2.25 ohms

L = 105 micro H

C = 72.5 micro F


a) f = 150 hz


capacitive reactance, XC = 1/(2*pi*f*C) = 1/(2*pi*150*72.5*10^-6) = 14.63 ohms

Inductive reactance, XL = 2*pi*f*L = 2*pi*150*105*10^-6 = 0.099 ohms

impedance, z = sqrt(R^2 + (XL - XC)^2)

= sqrt(2.25^2 + (14.63 - 0.099)^2)

= 14.7 ohms


b) f = 6.5 k hz = 6500 hz


capacitive reactance, XC = 1/(2*pi*f*C) = 1/(2*pi*6500*72.5*10^-6) = 0.338 ohms

Inductive reactance, XL = 2*pi*f*L = 2*pi*6500*105*10^-6 = 4.29 ohms

impedance, z = sqrt(R^2 + (XL - XC)^2)

= sqrt(2.25^2 + (0.338 - 4.29)^2)

= 4.55 ohms

c) Irms = Vrms/z = 5.66/14.7 = 0.385 A

d) Irms = Vrms/z = 5.66/4.55 = 1.24 A

e) fo = 1/(2*pi*sqrt(L*C))

= 1/(2*pi*sqrt(105*10^-6*72.5*10^-6))

= 1824 hz

f) Irms = Vrms/z

= Vrms/R

= 5.66/2.25

= 2.516 A